Izmir Construction is a company engaged in construction in Turkish west. On 1 January 2021 it issued 5,000 5-year bonds with a par value of $1,000 per bond. They have a current market price of $975, carry annual coupon rate of 9% and are callable at 105 anytime in 3rd, 4th or 5th year. The interest rate in year 3, 4 and 5 are 10%, 8% and 9%. Estimate the yield to call (YTC) and yield to maturity (YTM) and tell which rate is a better estimate of the expected rate of return on the bond


1
Expert's answer
2021-08-19T12:22:42-0400
"YTC=\frac{540}{(1+0.09)^1}+\frac{540}{(1+0.09)^2}+\frac{540}{(1+0.1)^3}+\frac{540}{(1+0.08)^4}+\frac{540}{(1+0.09)^5}=0.43298"

"YTM=\frac{540+\frac{1000-975}{5}}{\frac{1000+975}{2}}=0.55"